OpenStudy (anonymous):

Prove equation is an identity: csc(x)-sin(x)/csc(x) = cos^2(x)

OpenStudy (anonymous):

Should I substitute 1/sin(x) for csc? I'm not sure on how to simplify that and get 1-sin^2(x) so it equals cos^2(x)

OpenStudy (anonymous):

yes you should

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin (x) } - \sin (x)\div \frac{ 1 }{ \sin(x) } = \frac{1-\sin ^{2} }{ \sin (x)} \]

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

How do you get rid of the other sin(x)?

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (anonymous):

you're welcome! :) can you try looking at my ?

OpenStudy (anonymous):

yep